# Large Language Model Can Continue Evolving From Mistakes

### v3

@@ Abstract
As world knowledge evolves and new task paradigms emerge, Large Language Models (LLMs) often fall short of meeting new demands due to knowledge deficiencies and outdated information. Continual Learning (CL) is crucial for keeping LLMs up-to-date and addressing these deficiencies. However, traditional CL approaches struggle to balance task-width generality with task-depth specificity and often lack efficient data collection strategies, leading to increased training costs without addressing the model’s most critical needs.
Inspired by the ‘summarizing mistakes’ learning skill, we propose the Continue Evolving from Mistakes (CEM) method. This iterative approach continually evaluates LLMs to identify knowledge deficiencies based on their mistakes, collecting relevant data from multiple sources to supplement training in a targeted manner. To enhance the model’s utilization of supplemental knowledge and prevent forgetting, we developed three dataset construction strategies that integrate various types of continual pretraining (CPT) data and continual instruction tuning (CIT) data. Extensive experiments demonstrate the efficacy of the CEM method, achieving up to a 17% improvement in LLM accuracy in the best scenarios. Additionally, further experiments confirm the potential of combining CEM with other catastrophic forgetting mitigation strategies, enabling multi-round iterative optimization.
@@
[…]
@@ 1 Introduction
With the exponential growth of data and parameter sizes, coupled with the employment of instruction fine-tuning, Large Language Models (LLMs) have showcased superior performance across a wide array of downstream tasks Chang et al. (2023); Wei et al. (2022); Peng et al. (2023). However, in real-world applications, the ongoing emergence of new knowledge and task requirements necessitates continual task-specific adaptation and knowledge supplementation for LLMs. Otherwise, LLMs are likely to generate incorrect outputs based on outdated or lacking knowledge (i.e., ‘hallucinations’) Huang et al. (2023). Accordingly, Continual Learning(CL) is proposed to ensure LLMs’ alignment with the ever-changing world knowledge.
Given the significant training costs, it is more popular to continually pre-train or fine-tune LLMs on a sequence of emerging domain-specific or task-specific data, rather than periodically training them from scratch on the latest corpora (e.g., the most recently archived Wikipedia) Gururangan et al. (2020b). However, common CL tends to be task-width oriented, whereby the model is trained using new task-relevant data to enhance its applicability across a wider range of tasksGururangan et al. (2020b); Cossu et al. (2022), as opposed to focusing on the continual improvement of the model’s performance within a specific domain or task, i.e., task-depth oriented. In addition, building CL datasets is often challenging. Blind data collection fails to ensure access to the high-quality data that the model most requires for task-solving, and may introduce significant training costs due to irrelevant data.
To address these issues, we propose the Continue Evolving from Mistakes (CEM) method, aiming to provide a data-efficient method for continually enhancing the performance of LLMs on specific tasks through muiltiple evaluation and supplementation with mistake-relevant knowledge. Our method draws inspiration from the ‘summarizing mistakes’ learning skill of human students. For instance, when a student is struggling with math, they can leverage test mistakes to identify lacking knowledge and extract relevant information from the textbook. Through repeated math tests and summaries, improvements in their math scores have been noted. While recent studies have uncovered successful utilization of mistakes in constructing CoT dataAn et al. (2024), thereby enhancing the model’s performance in mathematical reasoning, there exists potential to extend the leveraging of mistakes to diverse tasks and integrate it within CL paradigms.
Recent studies have shown that the poor performance of LLM on specific tasks may result from two factors: (1) issues beyond its knowledge boundaries, indicating a lack of domain expertise Ren et al. (2023); Zhang et al. (2023b), and (2) knowledge errors within LLM, including those stemming from unreliable pre-training data sources and outdated information Rae et al. (2021); Zhao et al. (2023). We argue that mistakes effectively expose the LLM’s knowledge deficiencies and errors. Addressing these directly represents the most efficient approach to enhancing the model. Therefore, at the start of each CEM round, we test the model and gather cases of incorrect or poor performance. We deconstruct these ‘bad cases’ to identify the most lacking or erroneous knowledge points, then obtain supplemental data from search engines and online encyclopedias to address these gaps. To enhance the LLM’s knowledge acquisition from this supplemental data and prevent forgetting, we devised multiple dataset construction strategies utilizing both unsupervised Continual Pre-training (CPT) data and supervised Continual Instruction Tuning (CIT) dataWu et al. (2024). By integrating CEM with strategies to mitigate catastrophic forgetting, we enable iterative, continual model evolution.
@@
Our contributions are summarized as follows: (1) We propose the CEM method. This method explores knowledge boundaries from LLM’s mistakes, endowing LLMs with the capacity for continual iterative evolution. (2) To assist LLMs in acquiring and comprehending factual knowledge from supplemental corpora, and to mitigate knowledge forgetting, we propose three strategies for constructing supplemental training sets. (3) We conduct experiments on multiple open-source LLMs for two complex knowledge Question-Answering (QA) tasks. Across all scenarios, the CEM method consistently exhibit significant improvements, achieving an optimal accuracy enhancement of 17%. (4) We verify that CEM can be used together with methods for mitigating catastrophic forgetting Li et al. (2022), such as replay-based methods Chaudhry et al. (2019); Wu et al. (2021). This ensures that catastrophic forgetting can be suppressed even with multiple iterations.
@@ 2 Related Work
Instruction fine-tuning LMs often struggle to maintain alignment well with human intent after pre-training. However, this issue can be alleviated by employing techniques such as Instruction Fine-Tuning (IFT) Ouyang et al. (2022); Stiennon et al. (2022) and Reinforcement Learning from Human or AI Feedback (RLHF/RLAIF)Gao et al. (2022); Bai et al. (2022). IFT involves training an LM to generate responses based on provided prompts, which might optionally include task instructions. The data used for fine-tuning is usually less than the original pre-training corpus. These datasets used in IFT are often curated through crowd-sourcing or extracted from an LM that is already capable of generating instruction-following examplesHan et al. (2024). Studies have shown that IFT can enhance zero-shot generalization to various unseen tasks. To optimize the fine-tuning process, various strategies have been proposed to make it more parameter-efficient. These strategies include using adapters Hu et al. (2022), prompt tuning Li and Liang (2021), and other techniques.
@@
Continual pretraining involves regularly updating LLMs with the latest facts or adapting them to specialized domains. To continually pre-train LLMs to update facts, researchers utilize dynamic datasets to assimilate real-time data from various sources such as news feeds Sun et al. (2019), scholarly articles Cossu et al. (2022), and social media Cossu et al. (2022).
Continual pre-training enhances domain knowledge through two approaches: (1) domain-incremental pre-training, which accumulates knowledge across multiple domains, and (2) domain-specific continual learning, which refines a general model into a domain expert by training on domain-specific datasets and tasks. In domain-incremental pre-training, Cossu et al. (2022) continually pre-trained LLMs on new multimodal data streams, preparing them for various downstream tasks. Ke et al. (2023) introduces a soft-masking mechanism to update LLMs with domain corpora, aiming to boost performance while preserving general knowledge. Notably, there has been research on continuous pretraining for specific domains such as finance Xie et al. (2023) and e-commerce Ma et al. (2023).
@@
Continual Instruction Tuning involves continually fine-tuning LLMs to learn how to follow instructions and transfer knowledge for future tasks Zhang et al. (2023a). It can be broadly classified into two categories: 1) task-incremental CIT and 2) domain-incremental CIT.
Task-incremental CIT aims to continually fine-tune LLMs on a series of task-specific instructions to enable them to solve new tasks. A straightforward solution is to continuously generate instruction-tuning data for new tasks and directly fine-tune LLMs on it Wang et al. (2023). However, studies have shown this process would cause a catastrophic forgetting of the learned knowledge and problem-solving skills in previous tasks Kotha et al. (2024).
Domain-incremental CIT aims to continually fine-tuning of LLMs on domain-specific instructions to acquire the knowledge required for solving tasks in new domains. TAPT Gururangan et al. (2020a) adaptively tunes LLMs on domain-specific data from various domains, followed by evaluation of their text classification ability in each domain. AdaptLLM Cheng et al. (2024) enriches the raw training corpus with reading comprehension tasks, helping the model leverage domain-specific knowledge while enhancing prompting performance.
@@ 3 Methods
The objective of this work is to enhance the knowledge of LLM through continual learning, resulting in iterative improvements in model performance.
@@ 3.1 Task Formulation
Question-Answering (QA) tasks are commonly used to evaluate models’ world knowledge and reasoning skills. Our focus is on enhancing the models’ world knowledge rather than their reasoning abilities. These tasks usually involve a series of single-choice questions with multiple options. The ability of a LLM to correctly identify the correct answer within these questions indicates its possession of the relevant knowledge. Conversely, failure to do so suggests a deficiency in the model’s understanding of the knowledge pertinent to the question.
@@
Then, for each collected question $q_{i}$ that was answered incorrectly by the LLM, collect supplemental corpus $C_{S_{i}}$,
$C_{S_{i}}=Search(q_{i}),\quad C_{S}=\{C_{S_{1}},C_{S_{2}},C_{S_{3}},\dots\},$
@@ 3.2 CEM Method
We propose the CEM method, involving three main steps: (1) testing the LLM and collecting incorrectly answered questions, (2) collecting supplemental corpora from multiple data sources, and (3) constructing a supplemental training set and fine-tuning the LLM. Figure 1 shows the process.
@@
Testing LLM and Collecting Incorrectly Answered Questions. We utilize the LLM to answer questions from a QA dataset. Subsequently, by utilizing pattern matching to extract answers from the LLM’s responses, we verify the accuracy of these answers and collect the questions that the LLM has answered incorrectly.
Collecting Supplemental Corpora from Multiple Data Sources. We argue that the questions answered incorrectly by the LLM reveal knowledge deficiencies. Based on the implicit information contained within questions incorrectly answered by the LLM, we collect supplemental corpora from multiple data sources, including search engines and online encyclopedias.
(1) Search Engine: Search engines are characterized by fast response time, high relevance matching, and diverse and rich content. We search for questions that the LLM answers incorrectly. The search engine ranks results based on relevance. We select the top-k results with the highest relevance and access their detailed pages to crawl the main textual content as supplemental corpora. To prevent the LLM from simply memorizing answers, we filter out search results that may contain both the original question and its answer. Additionally, we optimize the strategy for crawling the corpus to filter out irrelevant information and spam, such as advertisements.
In subsequent experiments, we utilize Bing as the search engine and set $k=4$. Figure 4 in Appendix shows a case of corpus collection from Bing.
(2) Online Encyclopedia: Compared to using search engines for retrieval, online encyclopedias provide more credible and professional search results, despite having lower hit rates than search engines. Since the organization of encyclopedias typically relies on keyword entries, we conduct entity extraction from the questions and options of the collected incorrectly answered questions, using the extracted entity keywords for retrieval. The main text of the retrieved results serve as supplemental corpora.
In subsequent experiment, we use Wikipedia as the online encyclopedia and instruct GPT-3.5 using a few-shot prompt for entity extraction. Figure 5 in Appendix shows a case of corpus collection from Wikipedia.
Constructing Supplemental Training Set and Fine-Tuning LLM. The supplemental training sets are constructed separately based on the supplemental corpus collected from Wikipedia and Bing. They are then employed to fine-tune the LLM. The construction method of the training set will be explained in detail in Section 3.3.
After completing one round of supplemental training, the model is retested on the QA dataset to evaluate its performance and gather a new set of incorrectly answered questions. This can serve as the first step for the next round of supplemental training.
It is important to note that supplemental training should not be never-ending, as this may lead to severe catastrophic forgetting and the boundary effect of diminishing accuracy improvement. In the following experiments, we will examine how the number of supplemental training rounds affects model improvement.

### v4

@@ Abstract
As world knowledge evolves and new task paradigms emerge, Continual Learning (CL) is crucial for keeping Large Language Models (LLMs) up-to-date and addressing their shortcomings. In practical applications, LLMs often require both continual instruction tuning (CIT) and continual pre-training (CPT) to adapt to new task paradigms and acquire necessary knowledge for task-solving. However, it remains challenging to collect CPT data that addresses the knowledge deficiencies in models while maintaining adequate volume, and improving the efficiency of utilizing this data also presents significant difficulties. Inspired by the ’summarizing mistakes’ learning skill, we propose the Continue Evolving from Mistakes (CEM) method, aiming to provide a data-efficient approach for collecting CPT data and continually improving LLMs’ performance through iterative evaluation and supplementation with mistake-relevant knowledge. To efficiently utilize these CPT data and mitigate forgetting, we design a novel CL training set construction paradigm that integrates parallel CIT and CPT data. Extensive experiments demonstrate the efficacy of the CEM method, achieving up to a 17% improvement in accuracy in the best case. Furthermore, additional experiments confirm the potential of combining CEM with catastrophic forgetting mitigation methods, enabling iterative and continual model evolution.
@@
Haokun Zhao1, Haixia Han2, Jie Shi1, Chengyu Du 1, Jiaqing Liang3 thanks: Corresponding Author, Yanghua Xiao 1,2 Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University1 Shanghai Institute of Artificial Intelligence for Education, East China Normal University2 Shanghai Key Laboratory of Data Science, School of Data Science, Fudan University3 {hkzhao23,jshi22}@m.fudan.edu.cn, {haixiahan03,cydu2024}@gmail.com, {liangjiaqing,shawyh}@fudan.edu.cn
@@ 1 Introduction
With the exponential growth of data and parameter sizes, Large Language Models (LLMs) have demonstrated superior performance across numerous downstream tasks Chang et al. (2023); Wei et al. (2022); Peng et al. (2023). However, in real-world applications, the continual emergence of new knowledge and task requirements necessitates ongoing task adaptation and knowledge supplementation for LLMs. Otherwise, LLMs may fall into hallucinations and generate incorrect responses Huang et al. (2023). Therefore, Continual Learning (CL) is proposed to ensure LLMs’ alignment with the evolving world knowledge and demands.
Figure 1: The figure illustrates two potential triggers for poor model performance: (1) Task Paradigm Unfamiliarity, and (2) Lack of task-relevant knowledge. Unfamiliarity with the task paradigm can cause deviations from expected interaction styles, while insufficient task knowledge may lead to hallucinations. Instruction tuning addresses only the former.
Wu et al. (2024) categorize Continual Learning into Continual Pre-training (CPT) and Continual Instruction Tuning (CIT). CPT requires a larger volume of training data to enrich LLMs’ knowledge for new domains, while CIT continually fine-tunes LLMs on emerging task-specific instruction data to adapt the models to evolving task paradigms with much smaller training costs and data requirements Razdaibiedina et al. (2023); Gururangan et al. (2020b); Cossu et al. (2022). However, recent work has highlighted the limitations of instruction tuning, suggesting it is merely a superficial alignment process focused on learning interaction styles and LLMs primarily acquire knowledge during the pre-training stage Zhou et al. (2023). This will trigger the bad response shown in Figure 1, where LLMs’ poor performance may result from a lack of requisite knowledge to solve the task rather than unfamiliarity with the task paradigm Ren et al. (2023); Zhang et al. (2023b). CPT is more effective than CIT in addressing this issue.
However, using CPT to supplement the model with the lacking knowledge required for task-solving presents several issues: First, collecting CPT data is challenging, as it is difficult to a priori identify the lacking knowledge while ensuring a sufficiently large data scale. Blind data collection and insufficient data volume may introduce irrelevant information, weaken generalization, and pose overfitting risks Zhang et al. (2024); Gekhman et al. (2024). Second, due to the weak constraints of self-supervised CPT data during parameters updates and the scale gap compared to initial pre-training data, efficiently utilizing CPT data and avoiding catastrophic forgetting remain difficult.
To address these issues, recent work has explored converting CPT data into CIT data, such as transforming it into reading comprehension questions to avoid the problem of insufficient data volume. Cheng et al. (2024) or using highly relevant QA pairs for CIT before CPT to enhance the data utilization efficiency Jiang et al. (2024). However, these methods demand high-quality data and the assistance of oracle LLM, and they overlook the crucial challenge of CPT mentioned above. Inspired by the ‘summarizing mistakes’ learning skill of human students, we argue that a model’s mistakes in ’exams’ reflect its inherent knowledge deficiencies. CPT data that directly addresses these gaps is the most effective for supplementing the model’s knowledge. Currently, the successful utilization of mistakes in constructing CoT data An et al. (2024); Tong et al. (2024) has enhanced the model’s performance in mathematics and reasoning, indicating the potential to leverage mistakes in continual learning.
In this paper, we propose the Continue Evolving from Mistakes(CEM) method, aiming to provide a data-efficient method for collecting CPT data and continually enhancing the performance of LLMs through multiple evaluation and supplementation with mistake-relevant knowledge. Specifically, at the start of each CEM round, we test the model, identify cases of poor performance, and deconstruct them to pinpoint knowledge deficiencies. We then gather abundant supplemental data based on mistake analysis from search engines and online encyclopedias as CPT data to address these gaps. To efficiently utilize these CPT data and mitigate forgetting, we designed a novel CL training set construction paradigm that integrates parallel CIT and CPT data. By integrating CEM with methods to mitigate catastrophic forgetting (e.g., Random Replay) Li et al. (2022), we enable iterative and continual model evolution.
@@
Our contributions are summarized as follows: (1) We propose the CEM method, which identifies knowledge deficiencies from LLMs’ mistakes and collects abundant relevant CPT data, enabling continual iterative evolution. (2) To assist LLMs in acquiring and comprehending factual knowledge from CPT data and mitigating knowledge forgetting, we propose three strategies for constructing supplemental training set using a parallel paradigm with CIT and CPT data. (3) We conduct experiments on multiple open-source LLMs for two complex knowledge QA tasks. The CEM method consistently demonstrates significant improvements across all scenarios, , achieving up to a 17.00% accuracy enhancement. (4) We verify that CEM can be combined with methods to mitigate catastrophic forgetting, such as replay-based methods Chaudhry et al. (2019); Wu et al. (2021), ensuring suppression of catastrophic forgetting over multiple iterations.
@@ 2 Related Work
Continual pre-training involves regularly updating LLMs with the latest facts or adapting them to specialized domains Wu et al. (2024). To continually pre-train LLMs and update facts, researchers utilize dynamic datasets to assimilate real-time data from various sources such as news feeds Sun et al. (2019), scholarly articles Cossu et al. (2022), and social media Cossu et al. (2022).
Continual pre-training enhances domain knowledge through two approaches: (1) domain-incremental pre-training, which accumulates knowledge across multiple domains, and (2) domain-specific continual learning, which refines a general model into a domain expert by training on domain-specific datasets and tasks. In domain-incremental pre-training, Cossu et al. (2022) continually pre-trained LLMs on new multimodal data streams, preparing them for various downstream tasks. Ke et al. (2023) introduced a soft-masking mechanism to update LLMs with domain corpora, aiming to boost performance while preserving general knowledge. Notably, there has been research on continual pre-training for specific domains such as finance Xie et al. (2023) and e-commerce Ma et al. (2023).
@@
Continual Instruction Tuning involves continually fine-tuning LLMs to learn how to follow instructions and transfer knowledge for future tasks Zhang et al. (2023a). It can be broadly classified into two categories: (1) task-incremental CIT and (2) domain-incremental CIT Wu et al. (2024).
Task-incremental CIT aims to continually fine-tune LLMs on a series of task-specific instructions to enable them to solve new tasks. A straightforward solution is to continuously generate instruction-tuning data for new tasks and directly fine-tune LLMs on it Wang et al. (2023). However, studies have shown this process causes catastrophic forgetting of the learned knowledge and problem-solving skills from previous tasks Kotha et al. (2024).
Domain-incremental CIT aims to continually fine-tune LLMs on domain-specific instructions to acquire the knowledge required for solving tasks in new domains. TAPT Gururangan et al. (2020a) adaptively tunes LLMs on domain-specific data from various domains, followed by evaluation of their text classification ability in each domain. AdaptLLM Cheng et al. (2024) enriches the raw training corpus with reading comprehension tasks, helping the model leverage domain-specific knowledge while enhancing prompting performance.
@@ 3.1 Task Formulation
Question Answering (QA) tasks are commonly used to evaluate LLM’s domain knowledge. These tasks typically involve a series of single-choice questions with multiple options. If an LLM correctly answers these questions, it indicates a strong grasp of the relevant knowledge. Conversely, incorrect answers reveal deficiencies in understanding.
@@
Then, for each collected question $q_{i}$ that was answered incorrectly by the LLM, collect supplemental corpus $C_{S_{i}}$ as CPT data,
$C_{S_{i}}=Search(q_{i}),\quad C_{S}=\{C_{S_{1}},C_{S_{2}},C_{S_{3}},\ldots\},$
@@ 3.2 CEM Method
We propose the CEM method, as illustrated in Figure 2. The main steps are as follows:
@@
Step 1&2: Testing LLM and Collecting Incorrectly Answered Questions. We use the LLM to answer questions from a QA dataset. By employing pattern matching to extract answers from the LLM’s responses, we calculate the accuracy and collect the questions, along with their options, that the LLM answered incorrectly.
Step 3: Obtaining Supplemental Corpus from Multiple Data Sources. Based on the implicit information of knowledge deficiencies contained within questions incorrectly answered by the LLM, we collect supplemental corpus from multiple data sources, including search engines and online encyclopedias.
(1) Search Engine: Search engines are characterized by high relevance matching and diverse content. We search for questions the LLM answers incorrectly. The search engine ranks results by relevance, and we select the top-k results to access their detailed pages, extracting the main textual content as supplemental corpus. To prevent the LLM from merely memorizing answers, we filter out results that contain both the original question and its answer. Additionally, we optimize the corpus crawling strategy to exclude irrelevant information and spam, such as advertisements.
In subsequent experiments, we utilize Bing as the search engine and set $k=4$. Figure 6 in Appendix shows the process of corpus collection from Bing.
(2) Online Encyclopedia: Compared to search engines, online encyclopedias provide more credible and professional results, despite lower hit rates. Since encyclopedias are typically organized by keyword entries, we conduct entity extraction from the questions and options, using the extracted entity keywords for retrieval. The main text of the retrieved results serves as the supplemental corpus.
In subsequent experiment, we use Wikipedia as the online encyclopedia and instruct GPT-3.5 using a few-shot prompt for entity extraction. Figure 7 in Appendix shows the process of corpus collection from Wikipedia.
Step 4&5: Constructing Supplemental Training Set and Fine-Tuning LLM. The supplemental training sets are constructed separately based on the supplemental corpus collected from Wikipedia and Bing. They are then employed to fine-tune the LLM. The construction strategies of the training set will be explained in detail in Section 3.3.
After completing one round of supplemental training, the model is retested on the QA dataset to evaluate its performance and gather a new set of incorrectly answered questions. This can serve as the first step for the next round of CEM.
It should be noted that supplemental training should not be endless, as it may lead to catastrophic forgetting and diminishing returns in accuracy improvement. Subsequent experiments explore the impact of the number of training rounds on model performance.